OpenStudy (anonymous):

how would i got about solving this question? you toss a 52.1g ball straight up. The ball leaves your hand at 7.53m/s. How high above your hand does the ball go? Hint: The force of gravity does work.

OpenStudy (ujjwal):

V^2= U^2- 2gh put V=0 the value of h you get is your answer.

OpenStudy (anonymous):

where do i put the mass value?

OpenStudy (anonymous):

@ujjwal ?

OpenStudy (ujjwal):

nowhere.. that isn't even needed!

OpenStudy (anonymous):

i dont get the answer in my box, the anwer is 2.89m and thats not what im getting ....

OpenStudy (ujjwal):

you get 2.89.. use the formula correctly. put V=0, U=7.53 and g=9.8 .. you will then get a value 2.892903061.. (to be precise)

OpenStudy (anonymous):

nvm .. i figured out what i did wrong. thanks so much for the help :)